# A decade of the fast-varying ionospheric and magnetospheric magnetic fields

### v1

@@ Acknowledgements.
The input Swarm, CryoSat-2, and Grace-FO data are part of the Swarm product and are available at https://swarmhandbook.earth.esa.int/catalogue/. The input ground observatory data are available at. The CI core and lithospheric models subtracted from the data and the CI magnetospheric model used for comparison are available at https://swarmhandbook.earth.esa.int/catalogue/. The CHAOS-7.18 model used for comparison is available at https://spacecenter.dk/files/magnetic-models/CHAOS-7. Time series of estimated ionospheric and magnetospheric Gauss coefficients will be made available upon acceptance of the manuscript.
@@
Given only observatory data, a piece-wide constant field along with the observatory bias cannot be uniquely determined in a least squares sense. This has been noted already in Langel et al. (1982). Although not complicated, there is no reference to our knowledge where a clear mathematical proof or physical explanation is given. We thus document our mathematical proof here, followed by a physical explanation.
@@
In the $k$-th time bin, the data $\mathbf{d}_{k}\in\mathbb{R}^{N_{k}}$ measured at ground observatories are linearly related to the field coefficients $\mathbf{x}_{k}\in\mathbb{R}^{M_{0}}$ via the design matrix $\mathbf{G}_{k}\in\mathbb{R}^{N_{k}\times M_{0}}$, combined with a contribution from the observatory bias
$\mathbf{d}_{k}\simeq\mathbf{G}_{k}\mathbf{x}_{k}+\mathbf{b}_{k}.$ (23)
@@
Here $\mathbf{b}_{k}\in\mathbb{R}^{N_{k}}$ represents the observatory bias corresponding to the measurements in this time bin. We use $N_{k}$ and $M_{0}$ to denote the dimensionality of the data and field coefficients, respectively. The dimensionality of the field coefficients $M_{0}$ is assumed to be uniform for all time bins. Since the observatory biases are static biases, the bias term in the $k$-th time bin is simply sampled from the collective observatory biases according to the measurements available in this time bin, and hence can be related to the vector representing the collective observatory biases $\mathbf{b}\in\mathbb{R}^{M_{b}}$ by a sampling matrix $\mathbf{S}_{k}\in\mathbb{R}^{N_{k}\times M_{b}}$. Therefore, we have
$\mathbf{d}_{k}\simeq\mathbf{G}_{k}\mathbf{x}_{k}+\mathbf{S}_{k}\mathbf{b}=\begin{pmatrix}\mathbf{G}_{k}&\mathbf{S}_{k}\end{pmatrix}\begin{pmatrix}\mathbf{x}_{k}\\ \mathbf{b}\end{pmatrix}.$ (24)
@@
Since the bias $\mathbf{b}$ is shared across all time windows, one cannot solve each time window separately, but has to solve the full system, written as
$\mathbf{d}=\begin{pmatrix}\mathbf{d}_{1}\\ \vdots\\ \mathbf{d}_{K}\end{pmatrix}\simeq\begin{pmatrix}\mathbf{G}_{1}&&\mathbf{0}&\mathbf{S}_{1}\\ &\ddots&&\vdots\\ \mathbf{0}&&\mathbf{G}_{K}&\mathbf{S}_{K}\\ \end{pmatrix}\begin{pmatrix}\mathbf{x}_{1}\\ \vdots\\ \mathbf{x}_{K}\\ \mathbf{b}\end{pmatrix}=\widetilde{\mathbf{G}}\widetilde{\mathbf{x}}.$ (25)
@@
where $\widetilde{\mathbf{G}}\in\mathbb{R}^{N\times M}$ is the full design matrix for the problem, with $N=\sum_{k=1}^{K}N_{k}$ and $M=KM_{0}+M_{b}$. $\widetilde{\mathbf{x}}\in\mathbb{R}^{M}$ is the augmented model vector. Solving the least squares problem above is mathematically equivalent to solving the normal equation
$\widetilde{\mathbf{G}}^{\top}\widetilde{\mathbf{G}}\widetilde{\mathbf{x}}=\begin{pmatrix}\mathbf{G}_{1}^{\top}\mathbf{G}_{1}&&\mathbf{0}&\mathbf{G}_{1}^{\top}\mathbf{S}_{1}\\ &\ddots&&\vdots\\ \mathbf{0}&&\mathbf{G}_{K}^{\top}\mathbf{G}_{K}&\mathbf{G}_{K}^{\top}\mathbf{S}_{K}\\ \mathbf{S}_{1}^{\top}\mathbf{G}_{1}&\cdots&\mathbf{S}_{K}^{\top}\mathbf{G}_{K}&\sum_{k}\mathbf{S}_{k}^{\top}\mathbf{S}_{k}\end{pmatrix}\widetilde{\mathbf{x}}=\widetilde{\mathbf{G}}^{\top}\mathbf{d}.$ (26)
@@
We now want to calculate the rank or the nullity of the matrix $\widetilde{\mathbf{G}}^{\top}\widetilde{\mathbf{G}}\in\mathbb{R}^{M\times M}$. We assume that the submatrices $\mathbf{G}_{k}^{\top}\mathbf{G}_{k}$ are all invertible (i.e. the field coefficients can be estimated time-bin-wise if there is no bias on the measurements). In this case, we can apply the Guttman rank additivity formula
@@
$\mathrm{rank}\left(\widetilde{\mathbf{G}}^{\top}\widetilde{\mathbf{G}}\right)$ $=\mathrm{rank}\left(\mathbf{G}_{1}^{\top}\mathbf{G}_{1}\right)+\mathrm{rank}\begin{pmatrix}\mathbf{G}_{2}^{\top}\mathbf{G}_{2}&&\mathbf{0}&\mathbf{G}_{1}^{\top}\mathbf{S}_{1}\\ &\ddots&&\vdots\\ \mathbf{0}&&\mathbf{G}_{K}^{\top}\mathbf{G}_{K}&\mathbf{G}_{K}^{\top}\mathbf{S}_{K}\\ \mathbf{S}_{1}^{\top}\mathbf{G}_{1}&\cdots&\mathbf{S}_{K}^{\top}\mathbf{G}_{K}&\sum_{k}\mathbf{S}_{k}^{\top}\mathbf{S}_{k}-\mathbf{S}_{1}^{\top}\mathbf{G}_{1}\mathbf{G}_{1}^{+}\mathbf{S}_{1}\end{pmatrix}$ (28)
$=\sum_{k=1}^{K}\mathrm{rank}\left(\mathbf{G}_{k}^{\top}\mathbf{G}_{k}\right)+\mathrm{rank}\left(\sum_{k=1}^{K}\mathbf{S}_{k}^{\top}\mathbf{P}_{\mathbf{G}_{k}}^{\perp}\mathbf{S}_{k}\right)$
where $\mathbf{A}^{+}=(\mathbf{A}^{\top}\mathbf{A})^{-1}\mathbf{A}^{\top}$ is the Moore-Penrose pseudoinverse of matrix $\mathbf{A}$, and $\mathbf{P}_{\mathbf{A}}^{\perp}=\mathbf{I}-\mathbf{G}\mathbf{G}^{+}$ is the projection operator onto the orthogonal complement of the image of $\mathbf{A}$. Under our assumption, $\mathbf{G}_{k}$ is full column rank and $\mathrm{rank}(\mathbf{G}_{k}^{\top}\mathbf{G}_{k})=M_{0}$. Applying the rank-nullity theorem, we have
@@
$=\mathrm{dim}\left(\widetilde{\mathbf{G}}^{\top}\widetilde{\mathbf{G}}\right)-\sum_{k=1}^{K}\mathrm{rank}\left(\mathbf{G}_{k}^{\top}\mathbf{G}_{k}\right)-\mathrm{dim}\left(\sum_{k=1}^{K}\mathbf{S}_{k}^{\top}\mathbf{P}_{\mathbf{G}_{k}}^{\perp}\mathbf{S}_{k}\right)+\mathrm{dim}\,\mathrm{ker}\left(\sum_{k=1}^{K}\mathbf{S}_{k}^{\top}\mathbf{P}_{\mathbf{G}_{k}}^{\perp}\mathbf{S}_{k}\right)$
$=M-KM_{0}-M_{b}+\mathrm{dim}\,\mathrm{ker}\left(\sum_{k=1}^{K}\mathbf{S}_{k}^{\top}\mathbf{P}_{\mathbf{G}_{k}}^{\perp}\mathbf{S}_{k}\right)=\mathrm{dim}\,\mathrm{ker}\left(\sum_{k=1}^{K}\mathbf{S}_{k}^{\top}\mathbf{P}_{\mathbf{G}_{k}}^{\perp}\mathbf{S}_{k}\right)$
Therefore, the dimension of the nullspace of the original linear system can be identified with that of the matrix $\sum_{k}\mathbf{S}_{k}^{\top}\mathbf{P}_{\mathbf{G}_{k}}^{\perp}\mathbf{S}_{k}$. In a piece-wise constant field model with observatory data, not only does $\mathbf{S}_{k}$ sample the biases in each time bin, but the same sampling matrix also samples the design matrix. In other words, the design matrices in each time bin are sampled from the rows of a design matrix $\mathbf{G}_{0}\in\mathbb{R}^{M_{b}\times M_{0}}$, whose rows contain the collective design vectors for all observatories, hence $\mathbf{G}_{k}=\mathbf{S}_{k}\mathbf{G}_{0}$. The whole column space of $\mathbf{G}_{0}$ hence maps to zero under the matrix $\sum_{k}\mathbf{S}_{k}^{\top}\mathbf{P}_{\mathbf{G}_{k}}^{\perp}\mathbf{S}_{k}$
$\left(\sum_{k=1}^{K}\mathbf{S}_{k}^{\top}\mathbf{P}_{\mathbf{G}_{k}}^{\perp}\mathbf{S}_{k}\right)\mathbf{G}_{0}=\sum_{k=1}^{K}\mathbf{S}_{k}^{\top}\mathbf{P}_{\mathbf{G}_{k}}^{\perp}(\mathbf{S}_{k}\mathbf{G}_{0})=\sum_{k=1}^{K}\mathbf{S}_{k}^{\top}\mathbf{P}_{\mathbf{G}_{k}}^{\perp}\mathbf{G}_{k}=\mathbf{0}$ (30)
@@
$\mathrm{dim}\,\mathrm{ker}\left(\widetilde{\mathbf{G}}^{\top}\widetilde{\mathbf{G}}\right)=\mathrm{dim}\,\mathrm{ker}\left(\sum_{k=1}^{K}\mathbf{S}_{k}^{\top}\mathbf{P}_{\mathbf{G}_{k}}^{\perp}\mathbf{S}_{k}\right)\geq\mathrm{rank}(\mathbf{G}_{0})=M_{0}.$ (31)
@@
From a physical point of view, introduction of the observatory bias in the regular spherical harmonic analysis allows embedding an arbitrary time-invariant field that can be represented with the field coefficients in the bias. If a set of field coefficients solves the original least squares problem, adding an arbitrary constant vector $\delta\mathbf{x}$ to the field coefficients in all time bins uniformly also solves the problem, as long as the observatory biases are adjusted accordingly to compensate for the static field. Therefore there is no uniqueness in the solution.
@@
$\mathbf{d}_{k}=\begin{pmatrix}\mathbf{d}_{k,\mathrm{obs}}\\ \mathbf{d}_{k,\mathrm{sat}}\end{pmatrix},\quad\mathbf{G}_{k}=\begin{pmatrix}\mathbf{G}_{k,\mathrm{obs}}\\ \mathbf{G}_{k,\mathrm{sat}}\end{pmatrix},\quad\mathbf{S}_{k}=\begin{pmatrix}\mathbf{S}^{\prime}_{k}\\ \mathbf{0}\end{pmatrix}.$ (32)
@@
$\mathbf{x}_{k}=\begin{pmatrix}\mathbf{x}_{k}^{\mathrm{int}}\\ \mathbf{x}_{k}^{\mathrm{ion}}\\ \mathbf{x}_{k}^{\mathrm{mag}}\end{pmatrix},\quad\mathbf{G}_{k}=\begin{pmatrix}\mathbf{G}_{k,\mathrm{obs}}\\ \mathbf{G}_{k,\mathrm{sat}}\end{pmatrix}=\begin{pmatrix}\mathbf{G}_{k,\mathrm{obs}}^{\mathrm{int}}&\mathbf{G}_{k,\mathrm{obs}}^{\mathrm{ion}}&\mathbf{G}_{k,\mathrm{obs}}^{\mathrm{mag}}\\ \mathbf{G}_{k,\mathrm{sat}}^{\mathrm{int}}&\mathbf{G}_{k,\mathrm{sat}}^{\mathrm{ion}}&\mathbf{G}_{k,\mathrm{sat}}^{\mathrm{mag}}\end{pmatrix}.$ (33)
@@
For simplicity, we shall only discuss the case where the three fields are parametrised in the same set of spherical harmonics, hence $\mathbf{x}_{k}^{\mathrm{int}},\mathbf{x}_{k}^{\mathrm{ion}},\mathbf{x}_{k}^{\mathrm{mag}}\in\mathbb{R}^{M_{0}/3}$. A similar conclusion holds for a non-uniform truncation of the fields, but the notation will be much more complicated.
@@
Following eq. (13), we have $\mathbf{G}_{k,\mathrm{obs}}^{\mathrm{ion}}=\mathbf{G}_{k,\mathrm{obs}}^{\mathrm{mag}}=\mathbf{G}_{k,\mathrm{obs}}^{\mathrm{ext}}$, i.e. the design matrix linking the ionospheric and magnetospheric coefficients to the ground observations are the same, and are simply the design matrix for external coefficients $\mathbf{G}_{k,\mathrm{obs}}^{\mathrm{ext}}$ in conventional internal-external field separation. These sub-matrices corresponding to the ground observations are also sampled by the sampling matrix from the conglomerate design matrices, i.e.
$\mathbf{G}_{k,\mathrm{obs}}^{\mathrm{int}}=\mathbf{S}_{k}^{\prime}\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{int}},\quad\mathbf{G}_{k,\mathrm{obs}}^{\mathrm{ext}}=\mathbf{S}_{k}^{\prime}\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{ext}}.$ (34)
@@
On the other hand, following eq. (14), we have $\mathbf{G}_{k,\mathrm{sat}}^{\mathrm{mag}}=\mathbf{G}_{k,\mathrm{sat}}^{\mathrm{ext}}$, but $\mathbf{G}_{k,\mathrm{sat}}^{\mathrm{ion}}=\mathbf{G}_{k,\mathrm{sat}}^{\mathrm{int}}\bm{\Lambda}$. $\bm{\Lambda}$ is diagonal matrix responsible for changing the internal coefficients with ionospheric origins observed at satellite altitudes into the external coefficients with ionospheric origins observed on the ground, as described by eq. (12). In the end, the sub-matrix $\mathbf{G}_{k}$ takes the form
$\mathbf{G}_{k}=\begin{pmatrix}\mathbf{S}_{k}^{\prime}\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{int}}&\mathbf{S}_{k}^{\prime}\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{ext}}&\mathbf{S}_{k}^{\prime}\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{ext}}\\ \mathbf{G}_{k,\mathrm{sat}}^{\mathrm{int}}&\mathbf{G}_{k,\mathrm{sat}}^{\mathrm{int}}\bm{\Lambda}&\mathbf{G}_{k,\mathrm{sat}}^{\mathrm{ext}}\end{pmatrix}.$ (35)

### v2

@@ Acknowledgements.
The input Swarm, CryoSat-2, and Grace-FO data are part of the Swarm product and are available at https://swarmhandbook.earth.esa.int/catalogue/. The input ground observatory data are from INTERMAGNET and WDC data, accessed via the VirES platform (https://vires.services/). The CI core and lithospheric models subtracted from the data and the CI magnetospheric model used for comparison are available at https://swarmhandbook.earth.esa.int/catalogue/. The CHAOS-7.18 model used for comparison is available at https://spacecenter.dk/files/magnetic-models/CHAOS-7. Time series of estimated ionospheric and magnetospheric Gauss coefficients are available on Zenodo, at https://dx.doi.org/10.5281/zenodo.14787587.
@@
Given only observatory data, a piece-wise constant field along with the observatory bias cannot be uniquely determined in the least squares sense. This has been noted already in Langel et al. (1982). Although not complicated, there is no reference to our knowledge where a clear mathematical proof or physical explanation is given. We thus document our mathematical proof here, followed by a physical explanation.
@@
In the $i$-th time bin ($i=1,2,\cdots N_{t}$ where $N_{t}$ is the total number of time bins), the data $\mathbf{d}_{i}\in\mathbb{R}^{N_{i}}$ measured at ground observatories are linearly related to the field coefficients $\mathbf{x}_{i}\in\mathbb{R}^{M_{0}}$ via the design matrix $\mathbf{G}_{i}\in\mathbb{R}^{N_{i}\times M_{0}}$, combined with a contribution from the observatory bias
$\mathbf{d}_{i}\simeq\mathbf{G}_{i}\mathbf{x}_{i}+\mathbf{b}_{i}.$ (25)
@@
Here $\mathbf{b}_{i}\in\mathbb{R}^{N_{i}}$ represents the observatory bias corresponding to the measurements in this time bin. We use $N_{i}$ and $M_{0}$ to denote the dimensionality of the data and field coefficients, respectively. The dimensionality of the field coefficients $M_{0}$ is assumed to be uniform for all time bins. Since the observatory biases are static biases, the bias term in the $i$-th time bin is simply sampled from the collective observatory biases according to the measurements available in this time bin, and hence can be related to the vector representing the collective observatory biases $\mathbf{b}\in\mathbb{R}^{M_{b}}$ by a sampling matrix $\mathbf{S}_{i}\in\mathbb{R}^{N_{i}\times M_{b}}$. Therefore, we have
$\mathbf{d}_{i}\simeq\mathbf{G}_{i}\mathbf{x}_{i}+\mathbf{S}_{i}\mathbf{b}=\begin{pmatrix}\mathbf{G}_{i}&\mathbf{S}_{i}\end{pmatrix}\begin{pmatrix}\mathbf{x}_{i}\\ \mathbf{b}\end{pmatrix}.$ (26)
@@
Since the bias $\mathbf{b}$ is shared across all time bins, one cannot solve each time window separately, but has to solve the full system, written as
$\mathbf{d}=\begin{pmatrix}\mathbf{d}_{1}\\ \vdots\\ \mathbf{d}_{N_{t}}\end{pmatrix}\simeq\begin{pmatrix}\mathbf{G}_{1}&&\mathbf{0}&\mathbf{S}_{1}\\ &\ddots&&\vdots\\ \mathbf{0}&&\mathbf{G}_{N_{t}}&\mathbf{S}_{N_{t}}\\ \end{pmatrix}\begin{pmatrix}\mathbf{x}_{1}\\ \vdots\\ \mathbf{x}_{N_{t}}\\ \mathbf{b}\end{pmatrix}=\widetilde{\mathbf{G}}\widetilde{\mathbf{x}}.$ (27)
@@
where $\widetilde{\mathbf{G}}\in\mathbb{R}^{N\times M}$ is the full design matrix for the problem, with $N=\sum_{i=1}^{N_{t}}N_{i}$ and $M={N_{t}}M_{0}+M_{b}$. $\widetilde{\mathbf{x}}\in\mathbb{R}^{M}$ is the augmented model vector. Solving the least squares problem above is mathematically equivalent to solving the normal equation
$\widetilde{\mathbf{G}}^{\top}\widetilde{\mathbf{G}}\widetilde{\mathbf{x}}=\begin{pmatrix}\mathbf{G}_{1}^{\top}\mathbf{G}_{1}&&\mathbf{0}&\mathbf{G}_{1}^{\top}\mathbf{S}_{1}\\ &\ddots&&\vdots\\ \mathbf{0}&&\mathbf{G}_{N_{t}}^{\top}\mathbf{G}_{N_{t}}&\mathbf{G}_{N_{t}}^{\top}\mathbf{S}_{N_{t}}\\ \mathbf{S}_{1}^{\top}\mathbf{G}_{1}&\cdots&\mathbf{S}_{N_{t}}^{\top}\mathbf{G}_{N_{t}}&\sum_{i}\mathbf{S}_{i}^{\top}\mathbf{S}_{i}\end{pmatrix}\widetilde{\mathbf{x}}=\widetilde{\mathbf{G}}^{\top}\mathbf{d}.$ (28)
@@
We now want to calculate the rank or the nullity of the matrix $\widetilde{\mathbf{G}}^{\top}\widetilde{\mathbf{G}}\in\mathbb{R}^{M\times M}$. We assume that the submatrices $\mathbf{G}_{i}^{\top}\mathbf{G}_{i}$ are all invertible (i.e. the field coefficients can be estimated time-bin-wise if there is no bias on the measurements). In this case, we can apply the Guttman rank additivity formula
@@
$\mathrm{rank}\left(\widetilde{\mathbf{G}}^{\top}\widetilde{\mathbf{G}}\right)$ $=\mathrm{rank}\left(\mathbf{G}_{1}^{\top}\mathbf{G}_{1}\right)+\mathrm{rank}\begin{pmatrix}\mathbf{G}_{2}^{\top}\mathbf{G}_{2}&&\mathbf{0}&\mathbf{G}_{2}^{\top}\mathbf{S}_{2}\\ &\ddots&&\vdots\\ \mathbf{0}&&\mathbf{G}_{N_{t}}^{\top}\mathbf{G}_{N_{t}}&\mathbf{G}_{N_{t}}^{\top}\mathbf{S}_{N_{t}}\\ \mathbf{S}_{2}^{\top}\mathbf{G}_{2}&\cdots&\mathbf{S}_{N_{t}}^{\top}\mathbf{G}_{N_{t}}&\sum_{i}\mathbf{S}_{i}^{\top}\mathbf{S}_{i}-\mathbf{S}_{1}^{\top}\mathbf{G}_{1}\mathbf{G}_{1}^{+}\mathbf{S}_{1}\end{pmatrix}$ (30)
$=\sum_{i=1}^{N_{t}}\mathrm{rank}\left(\mathbf{G}_{i}^{\top}\mathbf{G}_{i}\right)+\mathrm{rank}\left(\sum_{i=1}^{N_{t}}\mathbf{S}_{i}^{\top}\mathbf{P}_{\mathbf{G}_{i}}^{\perp}\mathbf{S}_{i}\right)$
where $\mathbf{A}^{+}=(\mathbf{A}^{\top}\mathbf{A})^{-1}\mathbf{A}^{\top}$ is the Moore-Penrose pseudoinverse of matrix $\mathbf{A}$, and $\mathbf{P}_{\mathbf{A}}^{\perp}=\mathbf{I}-\mathbf{G}\mathbf{G}^{+}$ is the projection operator onto the orthogonal complement of the image of $\mathbf{A}$. Under our assumption, $\mathbf{G}_{i}$ is full column rank and $\mathrm{rank}(\mathbf{G}_{i}^{\top}\mathbf{G}_{i})=M_{0}$. Applying the rank-nullity theorem, we have
@@
$=\mathrm{dim}\left(\widetilde{\mathbf{G}}^{\top}\widetilde{\mathbf{G}}\right)-\sum_{i=1}^{N_{t}}\mathrm{rank}\left(\mathbf{G}_{i}^{\top}\mathbf{G}_{i}\right)-\mathrm{dim}\left(\sum_{i=1}^{N_{t}}\mathbf{S}_{i}^{\top}\mathbf{P}_{\mathbf{G}_{i}}^{\perp}\mathbf{S}_{i}\right)+\mathrm{dim}\,\mathrm{ker}\left(\sum_{i=1}^{N_{t}}\mathbf{S}_{i}^{\top}\mathbf{P}_{\mathbf{G}_{i}}^{\perp}\mathbf{S}_{i}\right)$
$=M-{N_{t}}M_{0}-M_{b}+\mathrm{dim}\,\mathrm{ker}\left(\sum_{i=1}^{N_{t}}\mathbf{S}_{i}^{\top}\mathbf{P}_{\mathbf{G}_{i}}^{\perp}\mathbf{S}_{i}\right)=\mathrm{dim}\,\mathrm{ker}\left(\sum_{i=1}^{N_{t}}\mathbf{S}_{i}^{\top}\mathbf{P}_{\mathbf{G}_{i}}^{\perp}\mathbf{S}_{i}\right)$
Therefore, the dimension of the nullspace of the original linear system can be identified with that of the matrix $\sum_{i}\mathbf{S}_{i}^{\top}\mathbf{P}_{\mathbf{G}_{i}}^{\perp}\mathbf{S}_{i}$. In a piece-wise constant field model with observatory data, not only does $\mathbf{S}_{i}$ sample the biases in each time bin, but the same sampling matrix also samples the design matrix. In other words, the design matrices in each time bin are sampled from the rows of a design matrix $\mathbf{G}_{0}\in\mathbb{R}^{M_{b}\times M_{0}}$, whose rows contain the collective design vectors for all observatories, hence $\mathbf{G}_{i}=\mathbf{S}_{i}\mathbf{G}_{0}$. The whole column space of $\mathbf{G}_{0}$ hence maps to zero under the matrix $\sum_{i}\mathbf{S}_{i}^{\top}\mathbf{P}_{\mathbf{G}_{i}}^{\perp}\mathbf{S}_{i}$
$\left(\sum_{i=1}^{N_{t}}\mathbf{S}_{i}^{\top}\mathbf{P}_{\mathbf{G}_{i}}^{\perp}\mathbf{S}_{i}\right)\mathbf{G}_{0}=\sum_{i=1}^{N_{t}}\mathbf{S}_{i}^{\top}\mathbf{P}_{\mathbf{G}_{i}}^{\perp}(\mathbf{S}_{i}\mathbf{G}_{0})=\sum_{i=1}^{N_{t}}\mathbf{S}_{i}^{\top}\mathbf{P}_{\mathbf{G}_{i}}^{\perp}\mathbf{G}_{i}=\mathbf{0}$ (32)
@@
$\mathrm{dim}\,\mathrm{ker}\left(\widetilde{\mathbf{G}}^{\top}\widetilde{\mathbf{G}}\right)=\mathrm{dim}\,\mathrm{ker}\left(\sum_{i=1}^{N_{t}}\mathbf{S}_{i}^{\top}\mathbf{P}_{\mathbf{G}_{i}}^{\perp}\mathbf{S}_{i}\right)\geq\mathrm{rank}(\mathbf{G}_{0})=M_{0}.$ (33)
@@
From a physical point of view, introducing the observatory bias in the regular spherical harmonic analysis allows one to embed an arbitrary time-invariant field that can be represented with the field coefficients in the bias. If a set of field coefficients solves the original least squares problem, adding an arbitrary constant vector $\delta\mathbf{x}$ to the field coefficients in all time bins uniformly also solves the problem, as long as the observatory biases are adjusted accordingly to compensate for the static field. Therefore there is no uniqueness in the solution.
@@
$\mathbf{d}_{i}=\begin{pmatrix}\mathbf{d}_{i,\mathrm{obs}}\\ \mathbf{d}_{i,\mathrm{sat}}\end{pmatrix},\quad\mathbf{G}_{i}=\begin{pmatrix}\mathbf{G}_{i,\mathrm{obs}}\\ \mathbf{G}_{i,\mathrm{sat}}\end{pmatrix},\quad\mathbf{S}_{i}=\begin{pmatrix}\mathbf{S}^{\prime}_{i}\\ \mathbf{0}\end{pmatrix}.$ (34)
@@
$\mathbf{x}_{i}=\begin{pmatrix}\mathbf{x}_{i}^{\mathrm{int}}\\ \mathbf{x}_{i}^{\mathrm{ion}}\\ \mathbf{x}_{i}^{\mathrm{mag}}\end{pmatrix},\quad\mathbf{G}_{i}=\begin{pmatrix}\mathbf{G}_{i,\mathrm{obs}}\\ \mathbf{G}_{i,\mathrm{sat}}\end{pmatrix}=\begin{pmatrix}\mathbf{G}_{i,\mathrm{obs}}^{\mathrm{int}}&\mathbf{G}_{i,\mathrm{obs}}^{\mathrm{ion}}&\mathbf{G}_{i,\mathrm{obs}}^{\mathrm{mag}}\\ \mathbf{G}_{i,\mathrm{sat}}^{\mathrm{int}}&\mathbf{G}_{i,\mathrm{sat}}^{\mathrm{ion}}&\mathbf{G}_{i,\mathrm{sat}}^{\mathrm{mag}}\end{pmatrix}.$ (35)
@@
For simplicity, we shall only discuss the case where the three fields are parametrised in the same set of spherical harmonics, hence $\mathbf{x}_{i}^{\mathrm{int}},\mathbf{x}_{i}^{\mathrm{ion}},\mathbf{x}_{i}^{\mathrm{mag}}\in\mathbb{R}^{M_{0}/3}$. A similar conclusion holds for a non-uniform truncation of the fields, but the notation will be much more complicated.
@@
Following Eq. (13), we have $\mathbf{G}_{i,\mathrm{obs}}^{\mathrm{ion}}=\mathbf{G}_{i,\mathrm{obs}}^{\mathrm{mag}}=\mathbf{G}_{i,\mathrm{obs}}^{\mathrm{ext}}$, i.e. the design matrix linking the ionospheric and magnetospheric coefficients to the ground observations are the same, and are simply the design matrix for external coefficients $\mathbf{G}_{i,\mathrm{obs}}^{\mathrm{ext}}$ in conventional internal-external field separation. These sub-matrices corresponding to the ground observations are also sampled by the sampling matrix from the conglomerate design matrices, i.e.
$\mathbf{G}_{i,\mathrm{obs}}^{\mathrm{int}}=\mathbf{S}_{i}^{\prime}\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{int}},\quad\mathbf{G}_{i,\mathrm{obs}}^{\mathrm{ext}}=\mathbf{S}_{i}^{\prime}\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{ext}}.$ (36)
@@
On the other hand, following Eq. (14), we have $\mathbf{G}_{i,\mathrm{sat}}^{\mathrm{mag}}=\mathbf{G}_{i,\mathrm{sat}}^{\mathrm{ext}}$, but $\mathbf{G}_{i,\mathrm{sat}}^{\mathrm{ion}}=\mathbf{G}_{i,\mathrm{sat}}^{\mathrm{int}}\bm{\Lambda}$. $\bm{\Lambda}$ is diagonal matrix responsible for changing the internal coefficients with ionospheric origins observed at satellite altitudes into the external coefficients with ionospheric origins observed on the ground, as described by Eq. (12). In the end, the sub-matrix $\mathbf{G}_{i}$ takes the form
$\mathbf{G}_{i}=\begin{pmatrix}\mathbf{S}_{i}^{\prime}\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{int}}&\mathbf{S}_{i}^{\prime}\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{ext}}&\mathbf{S}_{i}^{\prime}\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{ext}}\\ \mathbf{G}_{k,\mathrm{sat}}^{\mathrm{int}}&\mathbf{G}_{k,\mathrm{sat}}^{\mathrm{int}}\bm{\Lambda}&\mathbf{G}_{k,\mathrm{sat}}^{\mathrm{ext}}\end{pmatrix}.$ (37)